Shift Summary for Owl shift, 28/Sep/99

Miguel Barrio (mbarrio@hep.uchicago.edu)
Tue, 28 Sep 1999 07:28:21 -0500

Shift Summary for Owl shift, 28/Sep/99

On Shift: P. Shanahan, M. Barrio

Hours of beam available at all intensities: 6:30
Hours of beam available at nominal intensity: 6:30
Hours of E799 data (special runs not included): 6:00
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:
Many Kfrends needed to fix "Missing FFFF word" problem in crate 53.
Some pipeline resets as well. CsI ch. 1533 problematic.

========================================================================
Runs ended this shift:
Run 14749 - 14.0 M events - BEAM_799

Tapes brought to FCC this shift:
none.

========================================================================
Accesses during shift:
none.

========================================================================
Overview of shift:
E799 datataking

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14749 113 0:20 Laser RMS and Stability plots are way off.
RMS returns to normal next spill.

1:10 ONSPLIT requests tape UPE002 in ktev2's
tps3d2 drive to be mounted.

RF alarm in SDAQ terminal comes on sporadically,
clears after a while.

CsI timing: looks off by about 1-3ns (late).
This could be due to inherent laser timing and
to the recent RF shift, but the magnitude of
the effect in the laser is more than what these
explanations could account for. Check logbook
#22, page 7.

1:50 Intensity: > 9E12, dead time is about 5 secs
every other spill due to memory filling up.

2:00 Checking histos. Banana plots look veryu weird,
with a non-symmetric rejection. Check logbook
#22, page 8.

282 4:10 CsI ch. 1381 started to read out crap data,
bad capid, wacky laser response, etc. Reset
pipeline, kfrend-ed generously, problem fixed.

323-30 5:10 CsI ch. 1533 goes to -13% gain drift and stays
there. Reset pipeline with DBC to no avail.

363 5:55 CsI ch. 1533 still bad. E/p seems top track
laser gain drift. Measured HV looks fine and
agrees with DB,

14749 End of this run.

6:35 ArunT.: Effective Run 14750 (tapes UP*043-45;
UPS015), the definitions of outboxes 1 and 3
were changed for tape balancing.
Triggers: 7, 9, 13 and 15 were moved from
outbox 3 (drive 4, UPC tapes) to outbox 1
(drive 1, UPA tapes). The three tapes should
fill up evenly now, roughly anyway.

6:55 Dart-fresh-started to make changes above
effective.

14750 7:00 Started cosmic run. Will have an access to
check/fix CsI ch. 1533. Hogan and BobT are.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift